OpenStudy (ttop0816):

1+tan^2x/tan^2x would it be sec^2x csc^2x <<

OpenStudy (anonymous):

is it \(\frac{1+\tan^2(x)}{\tan^2(x)}\) or is it \(1+\frac{\tan^2(x)}{\tan^2(x)}\) ?

OpenStudy (ttop0816):

the first one!

OpenStudy (***[isuru]***):

1 + tan^2 x = sec^2 x sec^2 x = 1/ cos^2 x \[\frac{ \frac{ 1 }{ \cos^{2}x } }{ \frac{ \sin^{2} x }{ \cos^{2}x } }\]\[\frac{ 1 }{ \cos^{2}x} \times \frac{ \cos^{2} }{ \sin^{2}x }\]\[\csc^{2}x\] so u r correct !